Mathbee CBSE Mathbee · CBSE Class 12 Maths
NCERT Class 12 · Ch 13 Miscellaneous Exercise · Q9

NCERT Class 12 Maths · Chapter 13 Probability · Miscellaneous Exercise Question 9

Subsystem failure probabilities: P(A)=0.2, P(B alone)=0.15, P(A∩B)=0.15. Find (i) P(A|B), (ii) P(A alone).
Failure events: distinguish “alone” from total failure
Failure events: distinguish “alone” from total failure

💡 Concept Hints

  1. Define the events and identify which event is being conditioned on.
  2. Use P(E|F)=P(E∩F)/P(F), or separate the possible sources before applying total probability and Bayes.

✅ Step-by-Step Solution

Final Answer: (i) 1/2; (ii) 1/20.
  1. P(B)=P(B alone)+P(A∩B)=0.15+0.15=0.30.
  2. P(A|B)=0.15/0.30=1/2. P(A alone)=0.20−0.15=0.05=1/20.

Master Chapter 13: Probability

Practise interactive questions, flashcards, visual mind maps, and timed board tests for Probability.
📘 Chapter Notes & All Exercises 🚀 Open Interactive Workspace →
← Previous Question
Each entry of a 2×2 determinant is independently 0 or 1 with…
Next Question →
Bag I has 3 red,4 black; bag II has 4 red,5 black. Transfer a random…

All Questions in Probability — Miscellaneous Exercise

Related Questions in Probability